# Geometric theory of constrained Schrödinger dynamics with application to time-dependent density-functional theory on a finite lattice

### v1

@@ IX Conclusion
In this article, we have developed a geometric framework for Schrödinger dynamics with constraints and used this formalism to revisit the foundations of TDDFT. The resulting time-dependent Kohn–Sham-type equation involves a correction term that can either be interpreted as a local imaginary potential $w(t)$ or, better, a nonlocal exchange-type Hermitian operator (Eq. (82) or Eq. (85)). Numerical tests on the Hubbard dimer show that the geometric correction $w(t)$ has a very different structure from that of the standard exact time-dependent Kohn–Sham potential. This works paves the way for alternative approximations in TDDFT, which could potentially better describe systems in non-adiabatic regimes.
@@ Appendix C On the invertibility of $S^{\Psi}$
If $\gamma$ commutes with ${\mathds{1}}_{J}$ for a non trivial $J$, then ${\mathds{1}}_{J}$ is also in the kernel by (102). Hence $\ker(\widetilde{S}^{\gamma})$ has multiplicity 2 or more. Conversely, if $\gamma$ acts irreducibly, let us consider an arbitrary vector $v$ in the kernel of $\widetilde{S}^{\gamma}$, which means that $\gamma,v]=0$. Let $J$ be the set of indices $m$ so that $v_{m}=v_{1}$. Since $\gamma$ commutes with $v$ it must commute with the spectral projection ${\mathds{1}}_{J}={\mathds{1}}(v=v_{1})$ (by [72, Sec. 4.9]) and we conclude that necessarily ${\mathds{1}}_{J}\in\ker(\widetilde{S}^{\gamma})$. The irreducibility assumption tells us that $J=\{1,...,M\}$, hence $v$ is constant and the kernel has multiplicity one, as claimed. ∎

### v2

@@ IX Conclusion
In this article, we have developed a geometric framework for Schrödinger dynamics with constraints and used this formalism to revisit the foundations of TDDFT. The resulting time-dependent Kohn–Sham-type equation involves a correction term that can either be interpreted as a local imaginary potential $w(t)$ or, better, a nonlocal exchange-type Hermitian operator (Eq. (82) or Eq. (85)). Numerical tests on the Hubbard dimer show that the geometric correction $w(t)$ has a very different structure from that of the standard exact time-dependent Kohn–Sham potential. This works paves the way for alternative approximations in TDDFT, which could potentially better describe systems in non-adiabatic regimes. In [16] we study the continuous case and introduce a geometric Kohn–Sham scheme based on a universal functional (independent of the external potential).
@@ Appendix C On the invertibility of $S^{\Psi}$
If $\gamma$ commutes with ${\mathds{1}}_{J}$ for a non trivial $J$, then ${\mathds{1}}_{J}$ is also in the kernel by (102). Hence $\ker(\widetilde{S}^{\gamma})$ has multiplicity 2 or more. Conversely, if $\gamma$ acts irreducibly, let us consider an arbitrary vector $v$ in the kernel of $\widetilde{S}^{\gamma}$, which means that $\gamma,v]=0$. Let $J$ be the set of indices $m$ so that $v_{m}=v_{1}$. Since $\gamma$ commutes with $v$ it must commute with the spectral projection ${\mathds{1}}_{J}={\mathds{1}}(v=v_{1})$ (by [77, Sec. 4.9]) and we conclude that necessarily ${\mathds{1}}_{J}\in\ker(\widetilde{S}^{\gamma})$. The irreducibility assumption tells us that $J=\{1,...,M\}$, hence $v$ is constant and the kernel has multiplicity one, as claimed. ∎